OpenStudy (anonymous):

Last year a large trucking company delivered 8.0 x 10^5 tons of goods with an average value of $30,000 per ton. What was the total value of the goods delivered? Write the answer in scientific notation. 2.4 x 10^12 dollars 2.4 x 10^10 dollars 2.4 x 10^9 dollars 2.4 x 10^11 dollars

OpenStudy (anonymous):

Total value of delivered goods is: \[ 8.0 \times 10^5 \times 30000 = 2.4 \times 10^9 (dollars)\]